One way to maintain exclusivity for a brand is to raise its price.​ That's what luxury fashion and leather goods maker Louis Vuitton did. The company did not want the brand to become overexposed and too​ common, so it raised prices 20 percent and is slowing its expansion in China. The Louis Vuitton brand is the largest contributor to the​ company's $13.3 billion revenue from its fashion and leather​ division, accounting for​ $8 billion of those sales. It might seem counterintuitive to want to encourage fewer customers to purchase a​ company's products, but when price​ increases, so does the​ product's contribution​ margin, making each sale more profitable.​ Thus, sales can drop and the company can still maintain the same profitability as before the price hike. If a​ company's original contribution margin was 55​%, calculate the new contribution margin if price is increased 20​%.

Set the initial price equal to​ $1.00. Then the variable cost is ​$______.
Answer #1

Let the initial price be $1

Contribution margin = 55%

Hence, contribution margin = 1 x 55%

= $0.55

Variable cost per unit = Selling price per unit - Contribution margin per unit

= 1-0.55

= $0.45

Increase in price = 20%

= 1 x 20%

= 0.20

Contribution margin will also increase by = $0.20

Selling price per unit after increase = 1+0.20

= $1.20

Contribution margin per unit after increase = 0.55+0.20

= $0.75

New contribution margin ratio = Contribution margin per unit / Selling price per unit

= 0.75/1.20

= 62.5%

Kindly comment if you need further assistance.

Thanks‼!